# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

### v4

@@ V-E Scan matching observation
Table I: Results on NTU4DRadLM Dataset (Oculii Eagle). Best results in bold, second best results underlined. Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching, Abl.: Ablation. Relative translation errors ($t_{\mathrm{rel}}$) are expressed in %, relative rotation errors ($r_{\mathrm{rel}}$) in °/m, and time in ms
@@
NTU4DRadLM Baseline: GICP* [8] 4.13 0.0552 4.62 0.0184 4.84 0.0060 3.22 0.0060
NTU4DRadLM Baseline: Fast-LIO* [8] 2.94 0.0468 3.80 0.0208 7.16 0.0057 4.55 0.0064
4DRadarSLAM: APDGICP* [9] 3.56 0.0369 3.55 0.0171 6.09 0.0082 4.09 0.0097
@@
(Abl.) NDT Scan Matching [1] 12.61 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
(Abl.) GICP Scan Matching [2] 12.33 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
(Abl.) VGICP Scan Matching [15] 6.79 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
(Abl.) Gaussian Single Hypothesis 5.29 2.15 0.0318 2.46 0.0116 2.44 0.0030 3.97 0.0043
Gaussian Multi-Hypothesis 19.17 1.64 0.0310 2.41 0.0117 2.61 0.0036 3.71 0.0040
Table II: Results on Snail-Radar Dataset (Oculii Eagle and ARS548). Best results in bold, second best results underlined. Relative translation errors ($t_{\mathrm{rel}}$) are expressed in % and relative rotation errors ($r_{\mathrm{rel}}$) in mrad/m
@@
(Abl.) NDT Scan Matching [1] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
(Abl.) GICP Scan Matching [2] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
(Abl.) VGICP Scan Matching [15] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
(Abl.) Gaussian Single Hypothesis 3.36 0.5204 2.85 0.0689 2.84 0.0537 2.36 0.0927 4.00 0.1457
Gaussian Multi-Hypothesis 3.33 0.5147 2.45 0.0543 2.89 0.0498 2.55 0.1093 4.32 0.1569
@@
(Abl.) GICP Scan Matching [2] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
(Abl.) VGICP Scan Matching [15] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
(Abl.) Gaussian Single Hypothesis 1.34 0.3727 2.50 0.0487 2.77 0.0475 2.47 0.0826 2.37 0.0740
Gaussian Multi-Hypothesis 1.60 0.3573 4.02 0.0756 6.80 0.1000 2.93 0.0894 2.25 0.0746
Furthermore, we follow [31] and constrain the observation model described above by removing degrees of freedom that are deemed noisy in the scan matching due to the properties of the sensor. For example, radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms. In order to ignore the unreliable degrees of freedom, we transform the residual vector, observation matrix and covariance matrix by removing the rows and/or columns corresponding to Z/roll/pitch prior to applying the EKF update.
VI Experimental results
We use existing public 4D radar datasets for evaluation. In particular, we selected the NTU4DRadLM [8] and Snail-Radar [32] datasets, because together they cover multiple radar types (Oculii Eagle and ARS548), multiple robot platform types (handcart and car), and adverse conditions.
@@ VI Experimental results
Besides our Gaussian methods, we evaluate four ablations within our RIO framework: one without multi-hypothesis registration, and three using common established scan matching algorithms (NDT, GICP and VGICP) directly with radar point clouds. We use the current nominal state of the EKF directly as the initial estimate for single-hypothesis scan matching. The NDT version uses the implementation provided by the well known PCL (Point Cloud Library), while the GICP and VGICP versions use small_gicp [34] as the underlying implementation. In the case of Oculii Eagle scans, we set a target of 20 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian. In both cases we sample a swarm of 8 particles during scan matching, which we believe are enough for testing purposes without significantly increasing processing times. Other parameters include $t_{\mathrm{max}}$ = 15m and $\alpha_{\mathrm{max}}$ = 5° for keyframing, $d_{\mathrm{max}}$ = 4 for scan matching, and 0.5m/2° for multi-hypothesis particle dispersion. These parameters were chosen to suit the typical depth range/sparsity relation of a radar.
@@
VI-A Results on NTU4DRadLM
We evaluate using the NTU4DRadLM dataset [8], which includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU. Besides the evaluation performed with our pipeline, we also include previously published results of several radar-focused odometry methods, in order to provide context on the current performance of the state of the art. Due to a lack of Radar-Inertial methods, we also include radar-only methods. These include the GICP and Fast-LIO baselines tested in the original NTU4DRadLM paper, the APDGICP algorithm tested in the follow-up 4DRadarSLAM paper [9], as well as other works such as EFEAR-4D [10] or RIV-SLAM [11]. All works are categorized according to the sensory input used (IMU, egovelocity, scan matching). We include self-reported metrics as-is, as the trajectory files needed for metric calculation are usually not made public. The results for EKF-RIO [29] are generated by running the system directly, as there are no published results on this dataset. We report metrics without loop closure whenever possible, in order to focus on evaluating pure odometry.
@@ VI-A Results on NTU4DRadLM
Table I shows the comparison between all methods. Our Gaussian methods provide the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation). Moreover, in 3 out of the 4 sequences the multi-hypothesis scan matching approach provides improvements in translation error, especially in cp.
@@
Additionally, we report the average scan matching processing time for each method running under our odometry pipeline. The evaluation platform is a desktop PC equipped with an Intel Core i9-9900X CPU and 128 GB of RAM. The single-hypothesis version of our method achieves the fastest scan matching processing times, closely followed by VGICP. The multi-hypothesis approach, despite simultaneously registering 8 particles at once, is only 3.62x slower than the single-hypothesis version. We also mention that our current RIO implementation is entirely sequential, and that Gaussian modeling at keyframes consumes up to 608.78ms, raising the average scan processing time to 18.13ms. However, we believe that running the modeling and odometry tasks in parallel would remove any timing inconsistencies affecting real-time usage; which we leave as future work.
VI-B Results on Snail-Radar
@@
In addition to NTU4DRadLM, we perform additional testing on the Snail-Radar dataset [32], which uses both Oculii Eagle and ARS548 radars together with an Xsens MTi-3-DK IMU, and includes sequences recorded in harsher environment conditions. In particular, we selected the five sequences recorded on 2023/12/13 because of the nighttime and rain conditions of that day under a high speed car platform. The longest tested sequence (iaf_20231213_2) is 15 min in length, while the others are around 7 min.
@@
Table II shows results using the Oculii Eagle radar. In general, we can observe the best results to be spread across the different methods. Nonetheless, our Gaussian algorithms obtain the best or second best results in 9 out of 10 metrics, and the best translation error in 3 out of 5 sequences (the other two favoring NDT). The multi-hypothesis approach improves or matches results compared to the single hypothesis version in 3 out of 5 sequences. We also note that NDT seems to behave much better in Snail-Radar than in NTU4DRadLM.
@@
Table II also shows results using the ARS548 radar. This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points). We believe this to be the effect of Oculii Eagle’s scan enhancement algorithm. However, we can observe better metrics in general compared to Oculii Eagle. The sparser scans trigger a known issue in PCL’s implementation of NDT that cause a program termination, so unfortunately NDT is excluded from the comparison. The results once again show different methods leading in different sequences, however it is worth pointing out that 6 out of the 10 best results are now obtained by our Gaussian algorithms, as well as 3 out of the 10 second best results. Another thing to mention is that the multi-hypothesis scan matching seems to perform worse than the single-hypothesis version. We hypothesize this to be related to the increased sparsity of the scans, and believe it could be solved by improving the sampling distribution of the particles.
VI-C Additional qualitative experiments
In order to further justify our modeling approach based on globally optimized 3D Gaussians versus existing voxel-based approaches, we compare in Fig. 2 the output of our method against NDT. To do this, we prepare a radar-based map of NTU4DRadLM’s cp sequence using the supplied ground truth, and model a section of said map with both NDT and our method. The positions and sizes of the Gaussians generated by NDT are clearly constrained and influenced by the voxel grid, while our approach generates better fitted Gaussians. Moreover, our method is able to preserve geometric details such as corners, and allocates a higher budget of Gaussians to areas with higher point density, resulting in overall better representation of the geometry.
Additionally, we include a video supplement. In the video we show several Gaussian models of the same map created using different points-per-Gaussian ratios. It can be seen that, regardless of said hyperparameter, our modeling method is able to generate a representation reasonably fitting the map. Different number of Gaussians are allocated to areas of the map depending on the density and complexity of the area, as well as the available budget. Moreover, when the ratio is higher (meaning fewer Gaussians), sparse regions with few (outlier) radar points do not receive any Gaussians. Finally, we also include an example of the radar-inertial odometry pipeline on the same NTU4DRadLM cp sequence, compared to the ground truth trajectory.
VI-D Summary of results
@@
We consider that the results above validate our methodology. Our Gaussian modeling and scan matching algorithm produces comparable relative translation and rotation metrics to existing registration algorithms, as shown by the experiments. Moreover, in several sequences across several datasets and radar types our algorithm produces the best overall results. On the other hand, the multi-hypothesis approach seems to need further adjustments, although it is already capable of producing better results for some sequences.
@@
VII Conclusion
In this paper we introduce modeling and scan matching algorithms based on 3D Gaussians, which initialize and jointly optimize all parameters of a 3D Gaussian model to fit a radar point cloud, and make use of multiple hypothesis to support a scan matching process for subsequent incoming point clouds. Through experiments in multiple datasets with multiple radar types, we show that these methods produce results comparable to existing registration algorithms, and even manage to outperform them in several sequences. At the same time, our methods offer higher flexibility/expandability and faster scan matching compared to existing approaches based on individual points and voxelization [2, 1, 15, 16].
Future works include improving the run-time efficiency of the Gaussian modeling (replacing gradient descent, and running in parallel to the odometry pipeline), better initialization and dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as RCS (Radar Cross Section) values, and improving the robustness of the multi-hypothesis particle system with kinematic priors. Another line of work is further development of the RIO pipeline, using techniques such as Pose Graph Optimization (PGO), loop closure, and creating a global Gaussian map of the environment for the purposes of full SLAM.
Acknowledgment
This work was partially supported by the following grants: 1) INSERTION PID2021-127648OB-C31, and 2) NORDIC TED2021-132476B-I00 projects, funded by MCIN/AEI/ 10.13039/501100011033 and the “European Union NextGenerationEU / PRTR”.

### v5

@@ V-A Modeling experiments
Table I: Results on NTU4DRadLM Dataset (Oculii Eagle). Best results in bold, second best results underlined. Lower is better. Methods marked with * are self-reported metrics. EV: Egovelocity, SM: Scan Matching. Relative translation errors ($t_{\mathrm{rel}}$) are expressed in %, relative rotation errors ($r_{\mathrm{rel}}$) in °/m, and time in ms
@@
NTU4DRadLM baseline [8]: GICP* 4.13 0.0552 4.62 0.0184 4.84 0.0060 3.22 0.0060
NTU4DRadLM baseline [8]: Fast-LIO* 2.94 0.0468 3.80 0.0208 7.16 0.0057 4.55 0.0064
4DRadarSLAM [9]: APDGICP* 3.56 0.0369 3.55 0.0171 6.09 0.0082 4.09 0.0097
@@
NDT baseline [1] 12.61 2.42 0.0345 7.80 0.0408 5.61 0.0065 7.84 0.0100
GICP baseline [2] 12.33 2.93 0.0428 2.97 0.0111 2.99 0.0038 3.50 0.0042
VGICP baseline [15] 6.79 2.43 0.0362 2.56 0.0136 2.73 0.0036 3.84 0.0047
Gaussian (ours) ($K=1$) 5.29 2.26 0.0339 2.52 0.0117 2.26 0.0034 3.82 0.0043
Gaussian (ours) ($K=8$) 19.17 2.06 0.0266 2.71 0.0132 2.62 0.0039 3.41 0.0040
Gaussian (ours) ($K=128$) 433.52 2.31 0.0339 2.89 0.0138 2.67 0.0042 3.43 0.0039
Table II: Results on Snail-Radar Dataset (Oculii Eagle and ARS548). Lower is better. Best results in bold, second best results underlined. Relative translation errors ($t_{\mathrm{rel}}$) are expressed in % and relative rotation errors ($r_{\mathrm{rel}}$) in mrad/m
@@
NDT baseline [1] 4.13 0.5286 2.04 0.0519 3.95 0.0653 2.86 0.0786 3.15 0.0771
GICP baseline [2] 3.77 0.5118 2.93 0.0629 4.05 0.0629 2.77 0.1136 4.58 0.1572
VGICP baseline [15] 3.68 0.5141 3.06 0.0610 3.17 0.0551 2.53 0.1167 4.44 0.1489
Gaussian (ours) ($K=1$) 3.48 0.5112 2.97 0.0658 2.79 0.0537 2.48 0.1233 4.60 0.1564
Gaussian (ours) ($K=8$) 3.44 0.5165 4.15 0.0912 3.78 0.0676 2.33 0.0984 3.55 0.1202
Gaussian (ours) ($K=128$) 3.55 0.5507 3.43 0.0833 2.74 0.0480 2.34 0.1050 3.89 0.1227
@@
GICP baseline [2] 1.69 0.3049 3.36 0.0507 3.29 0.0487 2.42 0.0761 2.26 0.0796
VGICP baseline [15] 1.95 0.2798 2.57 0.0494 3.10 0.0491 2.81 0.0777 2.12 0.0836
Gaussian (ours) ($K=1$) 1.79 0.3561 3.37 0.0529 2.67 0.0380 2.37 0.0683 3.09 0.0902
Gaussian (ours) ($K=8$) 1.79 0.3561 3.56 0.0572 3.24 0.0424 2.14 0.0733 2.97 0.0883
Gaussian (ours) ($K=128$) 2.04 0.3580 3.67 0.0608 2.55 0.0420 2.29 0.0697 2.61 0.0986
Figure 3: Select trajectories of sequences from NTU4DRadLM and Snail-Radar, showing both the ground truth and various different scan matching algorithms used within our odometry pipeline.
Table III: Results of the ablation study on the modeling approach, replacing global/jointly optimized Gaussian model with local/voxelized VGICP model in our method. Lower is better. Best results in bold. Relative translation errors ($t_{\mathrm{rel}}$) expressed in %, relative rotation errors ($r_{\mathrm{rel}}$) in °/m (NTU4DRadLM) or mrad/m (Snail-Radar).
Radar/Sequence VGICP model Gaussian model
NTU4DRadLM
Oculii Eagle $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
cp 4.30 0.0651 2.26 0.0339
nyl 6.66 0.0333 2.52 0.0117
loop2 5.81 0.0067 2.26 0.0034
loop3 6.20 0.0081 3.82 0.0043
Avg. 5.74 0.0283 2.72 0.0133
Snail-Radar
Oculii Eagle $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
st_20231213_1 3.87 0.4654 3.48 0.5112
iaf_20231213_2 2.96 0.0696 2.97 0.0658
iaf_20231213_3 4.96 0.0783 2.79 0.0537
if_20231213_4 3.23 0.0845 2.48 0.1233
if_20231213_5 3.57 0.1112 4.60 0.1564
Avg. 3.72 0.1618 3.26 0.1821
ARS548 $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$ $t_{\mathrm{rel}}$ $r_{\mathrm{rel}}$
st_20231213_1 3.02 0.4107 1.79 0.3561
iaf_20231213_2 2.62 0.0492 3.37 0.0529
iaf_20231213_3 4.07 0.0418 2.67 0.0380
if_20231213_4 2.02 0.0805 2.37 0.0683
if_20231213_5 1.93 0.0825 3.09 0.0902
Avg. 2.73 0.1329 2.66 0.1211
V-B Scan matching sensitivity study
Afterwards, we individually test the stability and sensitivity of the Gaussian scan matching algorithm. For this purpose, we select the if_20231208_4 sequence from Snail-Radar due to its length (2.23 km over 8 min 35 s) and its challenging urban geometry. We extract radar scans (from both radars) every 30 s, which are then modeled with every algorithm (NDT, GICP, VGICP, Gaussians). We perform scan matching experiments between every model and 401 transformed copies of their source radar point clouds: 1 using the identity transformation, 100 using random translations, 100 random rotations, 100 random combined translations/rotations, and 100 copies of the scan with added noise. The random translations and rotations are uniformly sampled between 0 and 10 m/0°and 10°respectively. The direction of translation and axis of rotation is also uniformly sampled. The noise is normally distributed around zero with $\sigma=1$ m.
Figure 2 shows results for both Oculii Eagle and ARS548 radar scans. We can observe that our approach produces much lower translation and rotation errors across the board, with a much more compact error distribution. Failure rate is also improved compared to other algorithms in the case of random translations, combined transforms or noise. While it is higher in random isolated rotations, we believe this contributes to rejecting badly converged solutions and thus it can be said that our algorithm behaves more robustly. We must also mention that even though the NDT implementation is capable of reporting convergence failure this never seems to happen (failure rate of 0%), and thus the errors are much higher and spread out than for the other methods.
We report a single metric for our method in this experiment. This is because nearly all particles converge to the same solution, resulting in our algorithm obtaining the same metrics regardless of particle count. We think this happens because the same scan is used for both modeling and scan matching, and thus the likelihood of converging to a different local minimum is reduced. This also shows a strength of our algorithm: under “optimal” circumstances the aforementioned near full convergence of all particles occurs even though particle dispersion is high. The next experiment section is dedicated to a full odometry pipeline, where subsequent scans are registered against a model from a previous scan, which allows the multi-hypothesis strategy to take effect.
V-C Odometry experiments
Finally, we perform a comprehensive evaluation on a full Radar-Inertial Odometry task using both datasets and radar types. For this purpose we develop our own odometry pipeline following well established Extended Kalman Filter practice [31, 30, 12] to tightly integrate all incoming sensor information. Leveraging the Doppler information provided by 4D radar, scans are preprocessed with the RANSAC-LSQ algorithm [30] to both filter out outlier (dynamic) points, and estimate the current radar-frame velocity of the robot (egovelocity). The filter state includes position, velocity, attitude, accelerometer/gyroscope bias, and radar-to-IMU pose. We perform Kalman propagation using inertial data, and Kalman updates using radar egovelocity observations and scan matching results. In particular for the scan matching we follow [32] and constrain the observation model by only considering 3 degrees of freedom: X/Y/yaw. This is justified in the fact that radars have lower resolution in the Z axis, which degrades the usefulness of Z/roll/pitch information returned by scan matching algorithms.
@@ V-C Odometry experiments
V-C1 Results on NTU4DRadLM
We evaluate using the NTU4DRadLM dataset [8]. Besides the evaluation performed with our pipeline, we also include previously published results of several radar-focused odometry methods, in order to provide context on the current performance of the state of the art. Due to a lack of Radar-Inertial methods, we also include radar-only methods. These include the GICP and Fast-LIO baselines tested in the original NTU4DRadLM paper, the APDGICP algorithm tested in the follow-up 4DRadarSLAM paper [9], as well as other works such as EFEAR-4D [10] or RIV-SLAM [11]. All works are categorized according to the sensory input used (IMU, egovelocity, scan matching). We include self-reported metrics as-is, as the trajectory files needed for metric calculation are usually not made public. The results for EKF-RIO [30] are generated by running the system directly, as there are no published results on this dataset. We report metrics without loop closure whenever possible, in order to focus on evaluating pure odometry.
@@ V-C1 Results on NTU4DRadLM
Table I shows the comparison between all methods. Our Gaussian method provides the best or second best results for all sequences except for loop3, which seems to favor GICP-based methods (both baseline and our implementation). Performing multi-hypothesis scan matching helps in some sequences such as cp and loop3. Moreover, a high number of particles is not necessary to obtain these improvements.
@@
Additionally, we report the average scan matching processing time for each method running under our odometry pipeline. The evaluation platform is a desktop PC equipped with an Intel Core i9-9900X CPU and 128 GB of RAM. The single-hypothesis version of our method achieves the fastest scan matching processing times, closely followed by VGICP. When registering multiple particles our method shows non-linear scaling, as 8 particles only need 3.62x more time than a single particle, while 128 particles need 82x more time. We also mention that our current RIO implementation is entirely sequential, and that Gaussian modeling at keyframes consumes up to 608.78ms, raising the average scan processing time to 18.13ms. However, we believe that running the modeling and odometry tasks in parallel would remove any timing inconsistencies affecting real-time usage; which we leave as future work.
V-C2 Results on Snail-Radar
@@
In addition to NTU4DRadLM, we perform additional testing on the Snail-Radar dataset [28]. In particular, we selected the five sequences recorded on 2023/12/13 because of the nighttime and rain conditions of that day under a high speed car platform. The longest tested sequence (iaf_20231213_2) is 15 min in length, while the others are around 7 min.
@@
Table II shows results using the Oculii Eagle radar. In general, we can observe the best results to be spread across the different methods. Nonetheless, our Gaussian method obtains the best or second best results in 8 out of 10 metrics, and the best translation error in 3 out of 5 sequences (the other two favoring NDT). Our Gaussian approach produces the best translation errors in 3 out of 5 sequences, and the best rotation errors in 2 out of 5 sequences. The multi-hypothesis approach is capable of improving performance in 4 out of 5 sequences. Regarding the number of particles, a higher number benefits 3 out of 5 sequences, which indicates much more challenging geometry compared to NTU4DRadLM. We also note that NDT seems to behave much better in Snail-Radar than in NTU4DRadLM.
@@
Table II also shows results using the ARS548 radar. This radar produces much sparser point clouds than Oculii Eagle (around 10x fewer points). We believe this to be the effect of Oculii Eagle’s scan enhancement algorithm. However, we can observe better metrics in general compared to Oculii Eagle. The results once again show different methods leading in different sequences, however it is worth pointing out that our method produces the best results in 2 out of 5 sequences.
V-C3 Ablation study
In order to show the individual contribution of our global, jointly optimized Gaussian modeling approach, we perform an ablation study based on the odometry experiment combining our scan matching implementation with a strong locally constructed/voxelized baseline for modeling (VGICP). That is, we use the Gaussian models generated by VGICP with our scan matching approach and a single particle. Table III shows the results on all tested sequences. In the case of NTU4DRadLM our global models clearly obtain the best metrics on all sequences. For Snail-Radar the improvements are more marginal and reduced to some sequences, however our approach still obtains the best translation errors on average with both radars, while obtaining the best rotation errors on average with ARS548.
V-C4 Qualitative results
Fig. 3 shows plots of select sequences from both datasets. Specifically, we plot the loop2 sequence from NTU4DRadLM and the iaf_20231213_3 sequence from Snail-Radar, the latter for both Oculii Eagle and ARS548 radars. These two sequences are especially long and complex, which allows for qualitatively evaluating the performance of the different methods. We can observe that NDT drifts the most in loop2, while the rest stay fairly close to each other, and in particular our algorithm with 128 multi-hypothesis particles is the closest to the ground truth, closely followed by the 8 particle and 1 particle (single hypothesis) versions. As for iaf_20231213_3, the algorithms generally perform better with the ARS548. The 128 particle version of our algorithm is especially close to the ground truth using both radars.
V-D Summary of results
@@
We consider that the results above validate our methodology. Our Gaussian modeling and scan matching algorithm produces comparable or superior relative translation and rotation metrics to existing registration algorithms, as shown by the experiments. Moreover, in several sequences across several datasets and radar types our algorithm produces the best overall results. On the other hand, although the multi-hypothesis approach provides improvements in some sequences, it also produces regressions in others, indicating the need for further adjustments.
@@
VI Conclusion
In this paper we introduce a new modeling approach for radar point clouds based on globally and jointly optimized 3D Gaussians unlike existing voxelized or locally constructed approaches [1, 2, 15]. We show how this representation produces more robust and better performing scan matching, which we also propose to perform using multiple registration hypotheses. Through odometry experiments in multiple datasets with multiple radar types, we show that our approach produces results comparable to existing registration algorithms, managing to outperform them in several sequences. At the same time, our method offers higher expandability and faster scan matching compared to existing locally constructed or voxelized approaches [2, 1, 15, 16].
Future works include improving the run-time efficiency of the Gaussian modeling (replacing gradient descent, and running in parallel to the odometry pipeline), dynamic management of Gaussians (as in 3DGS), as well as incorporating radar-specific information such as RCS (Radar Cross Section) values, and improving the formulation of the multi-hypothesis particle system with kinematic priors.